write 6x^2-17x+13=20x^2-32 in standard form then solve.
Answer by mananth(16949)

write 6x^2-17x+13=20x^2-32 in standard form then solve.
...
6x^2-17x+13-20x^2+32=0
..
-14x^2-17x+45=0
...
-14x^2-35x+18x+45=0
-7x(2x+5)+9(2x+5)=0
(2x+5)(9-7x)=0
...
x=-5/2, x=9/7
...
m.ananth@hotmail.ca